OpenStudy (anonymous):

when he was 35 mr .ramos bought 300 shares of stock for $5 per share in a well-established company he received annual divdends in the amount of $1.25 per share.upon retiring at the age of 65 he sold his shares at a price of $68 per share. what was mr.Ramos return on investment.?

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

300 * 5 = 1500 30 * 300 * 1.25 = 300 *68 =

OpenStudy (nevermind_justschool):

Answer the rest of his formula and I or him will help you :)

OpenStudy (anonymous):

is it 20944?? @nevermind_justschool @Ziferot

OpenStudy (anonymous):

i don't know))) what did you do?

OpenStudy (anonymous):

I multiplied 300x68

OpenStudy (anonymous):

you must do everything of those)

OpenStudy (anonymous):

300 * 5 = 1500 30 * 300 * 1.25 = 300 *68 =

OpenStudy (anonymous):

300 *68 + 30 * 300 * 1.25 - 300 * 5=

OpenStudy (anonymous):

i got 38304750